The max no of different permutations of 4 letters of the word EARTHQUAKE

Grade:11

4 Answers

nath
7 years ago
The maximum number of different permutation of 4 letter of the word EARTHQUAKE is 44.
hope it will clear your question.
ALL THE BEST!
T Jahnavi
7 years ago
But the answer is 2190...
shubh
6 years ago
Case 1 4 distinct lettersPermutation = 8c4 × 4!Case 2 2 same 2 distinct lettersPermutation = 2c1 × 7c2 × 4!Case 32 same & 2 same lettersPermutation = 4! / 2!×2!Total no of ways = case 1+ case 2+ case 3Answer = 1680 + 510 + 6 = 2190
ARVIND GUPTA
6 years ago
There are 3 different casesCase 14 distinct letters Permutation = 8c4 × 4!Case 2 2 same 2 distinct letters Permutation = 2c1 × 7c2 × 4!Case 32 same & 2 same letters Permutation = 4! / 2!×2!Total no of ways = case 1+ case 2+ case 3Answer = 1680 + 514 + 6 = 2190